OpenStudy (anonymous):

bob has three times as many baseball cards as bill together they have 68 cards how many cards does bob have

OpenStudy (whpalmer4):

Write the equations representing what you know, then solve: Bob = 3 * Bill Bob + Bill = 68

OpenStudy (whpalmer4):

Bob = 68 - Bill but Bob also = 3 * Bill so 68 - Bill = 3 * Bill Add Bill to both sides and divide both sides by the right number to get Bill = ? Then plug that value into Bob + Bill = 68 to find out how many Bob has.

OpenStudy (anonymous):

i dont no how to solve it im so cionfused with math

OpenStudy (whpalmer4):

68 - Bill = 3 * Bill Add Bill to both sides (doing the same thing to both sides of the equation doesn't change the truth) 68 - Bill + Bill = Bill + 3*Bill 68 = 4*Bill Now divide both sides by 4 to find out the value of Bill 68/4 = Bill = ?

OpenStudy (anonymous):

new question raul is 5 years older than twice carlos age the sum of their age is 101 what is carlos age

OpenStudy (whpalmer4):

What did you get for your answers in the previous one? You should really post separate questions in separate posts. Again, write equations for what we know. Raul = Carlos * 2 + 5 Raul + Carlos = 101 Here we could substitute Carlos * 2 + 5 wherever we see Raul in the second equation: Carlos * 2 + 5 + Carlos = 101 Carlos * 3 + 5 = 101 Carlos * 3 + 5 - 5 = 101 - 5 Carlos * 3 = 96 Carlos = ?